Assume that personal happiness can be attributed to diet (73%) or non-dietary factors (27%). Non- dietary factors include exercise (63%) or weather objects (37%). Dietary factors are volume of water consumed (44%), average dark chocolate consumption (22%), and average number of egg rolls consumed (34%). a) b) Determine the probability that happiness is due to exercise. Determine the probability that happiness is due to dark chocolate or egg roll eating.
Answer #1

a)P(happiness is due to exercise)=P(non-dietary)*P(exercise |non-dietary)=0.27*0.63=0.1701

b)

P(happiness due to dark chocolate or egg roll)=P(dietary)*P(dark chocolate or egg roll|dietary)

=0.73*(0.22+0.34)=0.4088